# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 5 Stochastic PDR for quadratic problems (SPDRQ)
In this section, we will focus on the following stochastic PDRQ method
@@
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma Q+\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(N_{1}x^{k}+\sigma f-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right)\\ y^{k+1}=\left\{\begin{array}[]{ll}\tilde{y}^{k+1}_{i}=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{x+1}]&i=i_{k}\\ y^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}\tilde{y}_{\operatorname{test},i}^{k+1}=(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}_{i}^{k+1}-\bar{y}_{i}^{k}]&i=i_{k}\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k}\\ \end{array}\right.\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\tilde{\bar{y}}^{k+1}_{i}=\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k}\\ \bar{y}^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ \end{array}\right.$ (SPDRQ)
@@
where $N_{1}\geq 0$, $N_{2}\geq I$ and $N_{2}$ is separable, i.e., $N_{2}y=(N_{2,1}y_{1},\ldots,N_{2,n}y_{n})^{T}$. The preconditioned iteration (2.8) can be recovered from the update $x^{k+1}$ in the above (SPDRQ) with $N_{2}=I$ and $N_{1}$ as in (2.7). We define $x_{\operatorname{test}}^{k+1}=x^{k+1}$. Similarly, with $\mathcal{T}_{Q}$ defined in (2.10), we do one-step deterministic update based on the $u^{k}$ and get $\tilde{u}^{k+1}$, i.e.,
@@
where $x^{k+1}=\tilde{x}^{k+1}$ and $\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1}$. With Proposition 2 and Lemma 3, we can derive the following lemma and theorem similarly.
@@
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y))$ $\leq\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{Qp}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{Qp}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}}\Big{)}.$ (5.2)
@@
The iteration sequence $u^{k}=(x^{k},y^{k},\bar{y}^{k})$ produced by (SPDRQ) converges weakly to a fixed point of $\mathcal{T}_{Q}$, i.e., $u^{**}=(x^{**},y^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})\in\mathcal{Z}^{*}$ is one of the saddle points of the problem (1.1) with $Qx^{**}+\mathcal{K}^{*}y^{**}=f$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover the sequence $(x^{k+1}_{\operatorname{test}}$, $y^{k+1}_{\operatorname{test}})$ also converges weakly to the saddle point $(x^{**},y^{**})$ almost surely.
Proof.
Inserting $u^{*}=(x^{*},y^{*},\bar{y}^{*})$, a fixed point of $\mathcal{T}_{Q}$, in (5.2), we conclude that $(x^{*},y^{*})$ is a saddle point by Lemma 3. It follows that $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Denoting
$\Delta^{k+1}=\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{Qp}},$
we have
$\mathbb{E}^{k}\Delta^{k+1}\leq\Delta^{k}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}}=\Delta^{k}-\|\mathcal{T}_{Q}u^{k}-u^{k}\|^{2}_{\mathbf{M}_{Q}}.$ (5.3)
The remaining proof is similar to Theorem 1 and we omit it here. ∎
For the gap estimate, we give the following two lemmas which are similar to the SPDR case.
@@
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}<n(3C_{Q}+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C_{Q}$ is given in Lemma 10.
@@
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}<D_{Q}\Delta^{0}$, where
    $D_{Q}=3\max(\sigma(\|N_{1}\|\|A_{1}\|)^{2},\ \tau(\|(N_{2}-I)\|\|A_{2}\|)^{2},\ \tau\|A_{3}\|^{2})$
@@
    $M_{Q,K}=\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}((1+1/\underline{p})D_{Q}\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\underline{p}^{2}}^{2})+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
For the restricted gap estimate for (SPDRQ), we have the following theorem similarly.
@@
If $u=(x,y,\bar{y}),$ with $\bar{y}=y-\tau\mathcal{K}x$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ produced by (SPDRQ) converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K},$ (5.5)
where the constant $C_{\mathbb{B}}$ is given by,
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2}\Big{[}\|x^{0}-x\|_{\frac{N_{1}}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}\Big{]}$
$+\frac{1}{2}(\Delta^{0}+\frac{1}{\underline{p}}M_{Q,K})+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(n(3C_{Q}+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})$
$-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C_{Q}+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}].$ (5.6)
Here $C_{Q}$ is produced in Lemma 10, $M_{Q,K}$ is from Lemma 11, $\Delta^{0}=\|u^{0}-u^{*}\|_{\mathbf{M}_{Qp}}^{2}$, $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$ and $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{B}_{2}$ randomly. Then the same rate holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
@@ 6 Numerical Experiments
In our experiments, we focus on the following image deblurring problem (6.1) based on the Kullback-Leibler (KL) divergence and the anisotropic second-order Total Generalized Variation (TGV) [5] regularization
@@
Here $U$ and $V$ are finite-dimensional Hilbert spaces and $w=(w^{1},w^{2})$. The two-dimensional problem (6.1) is standard and all definitions along with discretizations will be made clear afterward. Let us first reformulate the problem (6.1) as the saddle-point problem. $\mathcal{G}_{1}^{*}$ is an appropriate data fidelity measure with noisy data. We employ the following KL divergence which is widely used in the presence of Poisson noise
@@
We also use the convention $00\log 0=$ 0 and $-a\log(0)=\infty$ for $a>0$. It can be readily checked that its Fenchel conjugate is
@@
where the second equation has to be understood in terms of the identification $W=U^{4}$. Consequently, the negative adjoint realizes a discrete negative divergence operator according to $\langle\mathcal{E}w,q\rangle_{W}=-\langle w,\operatorname{div}q\rangle_{V}$ for all $w\in V,q\in$ $W$. With $\Delta w=(\Delta w^{1},\Delta w^{2}),$ we can express the operator *$\mathcal{E}^{*}\mathcal{E}w=-\frac{1}{2}\Delta w-\frac{1}{2}\delta w$. Here $\operatorname{div}q$ and $\delta w$ are gives as
@@
6.1 The preconditioner
Setting $N_{2}=I$, our goal is to find efficient preconditioners to solve $Tx^{k+1}=b^{k}$, with $x^{k+1}=(u^{k+1},w^{k+1})$ and $b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k}=(b_{u}^{k},b_{w}^{k})$ and
@@ 6.1 The preconditioner
satisfies $T^{\prime}-T\geq 0$ [7]. It is clear that feasible preconditioners $M$ for $T^{\prime}$ are also feasible for $T$ (Since $M\geq T^{\prime}$, we have $M\geq T$). Consequently, we obtain
$x^{k+1}=x^{k}+M^{-1}\left(b^{k}+\left(T^{\prime}-T_{1}\right)x^{k}-T^{\prime}x^{k}\right),$ (6.9)
@@
$\left(b^{\prime}\right)^{k}$ $:=b^{k}+\left(T^{\prime}-T_{2}\right)x^{k}$ $=\left[\begin{array}[]{c}b_{u}^{k}+(\sigma\tau I-\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1})u^{k}\\ b_{w}^{k}+\frac{\sigma\tau}{2}\left(\delta-\Delta\right)w^{k}\end{array}\right].$ (6.10)
Then the block symmetric Gauss–Seidel preconditioner for $T^{\prime}$ as in [7] is employed.
6.2 Parameters

### v3

@@ 5 Stochastic PDR for quadratic problems (SRPDRQ)
In this section, we will focus on the following relaxed and stochastic PDRQ method directly for compactness
@@
$\left\{\begin{array}[]{ll}\tilde{x}^{k+1}=\left(N_{1}+\sigma Q+\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(N_{1}x^{k}+\sigma f-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right)\\ x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}\tilde{x}_{t}^{k+1},\\ y^{k+1}=\left\{\begin{array}[]{ll}(1-\rho_{k})y_{i}^{k}+\rho_{k}\tilde{y}^{k+1}_{t,i},\quad\tilde{y}^{k+1}_{t,i}:=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}\tilde{x}^{x+1}]&i=i_{k}\\ y^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}^{k+1}_{t,i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+\rho_{k,y}(y_{\operatorname{test},i}^{k+1}-\tilde{y}_{t,i}^{k+1})&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\\ \end{array}\right.$ (SRPDRQ)
where $N_{1}\succeq 0$, $N_{2}\succeq I$ and $N_{2}$ is separable, i.e., $N_{2}y=(N_{2,1}y_{1},\ldots,N_{2,n}y_{n})^{T}$. Here $\rho_{k}$, $\rho_{k,y}$ are non-decreasing and have lower and upper limit $\rho_{l},\rho_{u}$, and $20<\rho_{l}<\rho_{u}<$ 2 which is similar as the SRPDR. It can be observed that the updates for $\bar{x}^{k+1}$ are not required in the SRPDRQ iterations, unlike in SRPDR. As a result, SRPDRQ can be more efficient in terms of iteration time.
To clarify, we refer to the general SRPDRQ as the stochastic fully relaxed preconditioned Douglas-Rachford method for quadratic problems (f-SRPDRQ) when $\rho_{k}\neq 1$ and $\rho_{k,y}\neq 1$. While when $\rho_{k}=1$, and $\rho_{k,y}=1$, we call it the stochastic preconditioned Douglas-Rachford method for quadratic problems (SPDRQ).
@@
If we set $N_{2}=I$ and $N_{1}$ as in (2.11), we derive the preconditioned iteration (2.12). Define $x_{\operatorname{test}}^{k+1}=x^{k+1}$ and $\mathcal{T}_{Q}$ as in (2.14), we do one-step deterministic update based on the $u^{k}$ and get $\tilde{u}^{k+1}$, i.e.,
@@
where $x^{k+1}=\tilde{x}^{k+1}$ and $\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1}$. Noting $I_{q,\varrho_{k}}:=\textrm{Diag}[\rho_{k},\rho_{k},\rho_{k,y}]$ as before, with Proposition 3 and Lemma 3, we can derive the following lemma and theorem similarly.
@@
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y))$ $\leq\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{Qp}I_{q,\varrho_{k}}^{-1}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{Qp}I_{q,\varrho_{k}}^{-1}}$ (5.2)
$-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}(2I-I_{q,\varrho_{k}})I_{q,\varrho_{k}}^{-2}}\Big{)}.$
Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set of the fixed points for $\mathcal{T}_{Q}$.
@@
The iteration sequence $u^{k}=(x^{k},y^{k},\bar{y}^{k})$ produced by SRPDRQ converges weakly to an $\mathbb{F}$-valued random variable $u^{**}=(x^{**},y^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})$ is a $\mathcal{Z}^{*}$-valued random variable with $Qx^{**}+\mathcal{K}^{*}y^{**}=f$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover the sequence $(x^{k+1}_{\operatorname{test}}$, $y^{k+1}_{\operatorname{test}})$ also converges weakly to the $\mathcal{Z}^{*}$-valued random variable $(x^{**},y^{**})$ almost surely.
For the gap estimate, we give the following two lemmas which are similar to the SRPDR case.
@@
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}<\frac{(3C_{Q}+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C_{Q}$ is given in Lemma 10.
@@
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}<\frac{D_{Q}\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}$, where
    $D_{Q}=3\max(\sigma\|N_{1}\|\|A_{1}\|^{2},\ \tau\|N_{2}-I\|\|A_{2}\|^{2},\ \tau\|A_{3}\|^{2})$
@@
    $M_{Q,K}=\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}((1+1/\underline{p})\frac{D_{Q}\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\underline{p}^{2}}^{2})+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Concerning the restricted gap estimate for SRPDRQ, we have the following similar theorem.
@@
If $u=(x,y,\bar{y}),$ with $\bar{y}=y-\tau\mathcal{K}x$, let $u^{k}$ be the iteration sequence generated by SRPDRQ, and let $z^{k}_{\operatorname{test}}=(x^{k},y^{k}_{\operatorname{test}})$ be the corresponding transitional variables. Denote $K_{1}:=K+1$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}=\frac{1}{K_{1}}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K_{1}}=\mathcal{O}({1}/{K}),$ (5.4)
where the constant $C_{\mathbb{B}}$ is a positive constant depending on $u^{0}$, $\mathcal{K}$, $\mathbf{P}$, $\rho_{l}$, $\rho_{u}$, $\mathbf{M}_{Q_{p}}$, $u^{*}$, $z$, and $C_{Q}$, $M_{Q,K}$. Here $C_{Q}$ is produced from Lemma 10, and $M_{Q,K}$ is from Lemma 11. Therefore the same rate holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
Remark 2.
Theorem 2, 3, 5, and 6 can also cover the situation of random sets of indices to be selected and updated. Sample the dual variable $y=(y_{1},\dots,y_{n})$ properly [36, 37, 13] into multiple blocks $\mathbb{S}_{1}$, $\mathbb{S}_{2},\dots,\mathbb{S}_{b}$, such that $\mathbb{S}_{1}\cup\mathbb{S}_{2}\cup\dots\cup\mathbb{S}_{b}=\{y_{1},y_{2},\dots,y_{n}\}$ and $\mathbb{S}_{i}\cap\mathbb{S}_{j}=\emptyset$ for any $i,j\in\{1,\dots,b\}$, $i\neq j$. We denote the probability of selecting a block $i\in\{1,\ldots,b\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and $\sum_{i=1}^{b}p_{i}=1$. Following the same routine, we can derive the same results on almost sure convergence and sublinear convergence of restricted primal-dual gap.
@@ 6 Numerical Experiments
In this section, we will illustrate two experiments, one is a deblurring problem, and the other is a Binary classification problem with real data.
6.1 TGV-Deblur Problem
In the first experiments, we focus on the following image deblurring problem (6.1) based on the Kullback-Leibler (KL) divergence and the anisotropic second-order Total Generalized Variation (TGV) [6] regularization
@@ 6.1 TGV-Deblur Problem
A similar denoising problem in multi-modal electron tomography with KL divergence and TGV regularization is studied in [29]. Here $U$ and $V$ are finite-dimensional Hilbert spaces and $w=(w^{1},w^{2})$. The two-dimensional problem (6.1) is standard and all definitions along with discretizations will be made clear afterwards. Let us first reformulate the problem (6.1) as the saddle-point problem. $\mathcal{G}_{1}^{*}$ is the appropriate data term measure with blurred data given by the following KL divergence [29, Equation (9)], which is a variant of [28, Equation (13)].
@@
where $b_{i}$ is the blurred image and we use the convention $00\log 0=$ 0 and $-a\log(0)=\infty$ for $a>0$. It can be readily checked that its Fenchel conjugate is
@@
where the second equation has to be understood in terms of the identification $W=U^{4}$. Consequently, the negative adjoint realizes a discrete negative divergence operator according to $\langle\mathcal{E}w,q\rangle_{W}=-\langle w,\operatorname{div}q\rangle_{V}$ for all $w\in V,q\in$ $W$. With $\Delta w=(\Delta w^{1},\Delta w^{2}),$ we can express the operator $\mathcal{E}^{*}\mathcal{E}w=-\frac{1}{2}\Delta w-\frac{1}{2}\delta w$. Here $\operatorname{div}q$ and $\delta w$ are gives as
@@
6.1.1 The preconditioner
Setting $N_{2}=I$, our goal is to find efficient preconditioners to solve $Tx^{k+1}=b^{k}$, with $x^{k+1}=(u^{k+1},w^{k+1})^{T}$ and $b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k}=(b_{u}^{k},b_{w}^{k})^{T}$ and
@@ 6.1.1 The preconditioner
satisfies $T^{\prime}-T\succeq 0$ [9]. It is clear that feasible preconditioners $M$ for $T^{\prime}$ are also feasible for $T$ (Since $M\succeq T^{\prime}$, we have $M\succeq T$). Consequently, we obtain
$x^{k+1}=x^{k}+M^{-1}\left(b^{k}+\left(T^{\prime}-T\right)x^{k}-T^{\prime}x^{k}\right),$ (6.9)
@@
$\left(b^{\prime}\right)^{k}$ $:=b^{k}+\left(T^{\prime}-T\right)x^{k}$ $=\left[\begin{array}[]{c}b_{u}^{k}+(\sigma\tau I-\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1})u^{k}\\ b_{w}^{k}+\frac{\sigma\tau}{2}\left(\delta-\Delta\right)w^{k}\end{array}\right].$ (6.10)
Now solving $Tx^{k+1}=b^{k}$ by preconditioned iteration: $x^{k+1}=x^{k}+M^{-1}(b^{k}-Tx^{k})$ is transformed into a new preconditioned iteration: $x^{k+1}=x^{k}+M^{-1}((b^{\prime})^{k}-T^{\prime}x^{k})$, which is much easier to compute and implement. We refer to [9, Section 5.2] for the block symmetric Gauss–Seidel preconditioners.
6.1.2 Parameters